# Harnessing Structured Knowledge: A Concept Map-Based Approach for High-Quality Multiple Choice Question Generation with Effective Distractors

### v1

@@ 1 Introduction
Multiple choice questions (MCQs), a widely used objective assessment format, require learners to identify the correct answer from options that include plausible but incorrect alternatives, known as distractors [8]. MCQ-based testing serves two functions: assessment and learning enhancement. Studies have shown that the act of retrieving information while answering MCQs strengthens memory retention and deepens understanding in learners [25, 9]. Evidence from empirical studies demonstrates that MCQs accelerate learning outcomes. Their objective nature allows straightforward evaluation and rapid feedback, essential for continuous assessments in a lesson plan [39, 6]. When distractors are thoughtfully designed to mirror common cognitive errors, MCQs can uncover learning deficiencies and reduce random guessing [18]. This allows educators to modify teaching methods and offer targeted student support [11]. However, creating quality MCQs with effective distractors that challenge learners at different cognitive levels is resource intensive and requires significant domain expertise [1]. To bridge this gap, we propose a novel framework that utilizes concept maps and instruction-tuned LLMs for automated MCQ generation and assessment. The GitHub repository for the code and concept map is available as a supplementary resource for researchers interested in replicating or building upon our work. [^1] GitHub Repository
@@
Traditional MCQ generation tackled question stem generation, key answer identification, and distractor generation separately. The generation of question stems used rule-based transformations and templates, the identification of key answers relied on domain-specific keywords and phrase matching [12], and corpus-based methods focusing on linguistic features or knowledge-based methods for distractors [23]. These methods suffered from rigid question templates, a lack of contextual understanding in key answer identification, and difficulties in generating plausible, semantically relevant distractors, creating major bottlenecks in the creation of high-quality MCQs at scale [1]. Deep learning-based approaches, particularly sequence-to-sequence models [37] with attention mechanisms [4], offered improvements, particularly the hierarchical recurrent encoder-decoder (HRED) architecture [17] with dynamic and static attention [22]. Subsequent studies explored different attention mechanisms to enhance distractor generation, such as co-attention further [33] for passage-question interaction [47], topic-aware attention using Latent Dirichlet Allocation [34], and SoftSel to mitigate answer-revealing distractors [41].
@@
Large-scale datasets played a crucial role in improving MCQ and distractor generation. While datasets like RACE [44] and SWAG [44] focused on assessing reading comprehension and reasoning abilities, domain-specific datasets such as SciQ [43] and EduQG [19] improved structured question generation in different subjects. Distractor selection based on semantic similarity has been done employing cosine similarity of word vectors using these datasets, pre-trained language models (PLMs) such as word2vec [28], GloVe [29] and fastText [7]. Transformer-based models, such as GPT [30], BERT [13], T5 [31], and BART [21], further enhanced distractor generation through fine-tuning and retrieval-augmented pre-training approaches. Recent studies leverage in-context learning [27] and single-stage and multistage prompting techniques [14, 24] to improve distractor quality. These investigations have identified significant room for improvement in the generation of domain-specific and cognitively diverse questions with effective distractors. Furthermore, while promising, LLMs introduce new challenges such as bias [16] and hallucination [45], which could lead to seemingly accurate, yet factually inaccurate key answers and distractors. Although knowledge integration shows promise in grounding model output, current methods still struggle to fully address these challenges, necessitating expert validation, which creates significant resource bottlenecks.
@@ 2.1 Content and Concept Map Design
The concept map is the central piece of our automated MCQ generation framework. To systematically organize the content, we created a hierarchical concept map encompassing 19 units of classical physics using the OpenStax High School Physics textbook [40]. These units encompassed five major topics in physics (Figure 1(1a)): Mechanics, Electricity and Magnetism, Waves and Acoustics, Thermodynamics, and Optics. This hierarchical structure represents a structured knowledge repository to accommodate future expansion with additional subjects and grade levels.
@@
Within our concept map (Figure 1(1b)), each ‘unit’ is aligned with a unit from the OpenStax textbook, encapsulating ‘topics’ that are further divided into ‘subtopics’. ‘Topics’ align with sections in the OpenStax textbook, and for each topic, we mapped learning objectives to the cognitive and knowledge dimensions of the revised Bloom’s taxonomy [3]. This mapping ensures that the knowledge structure not only captures content but also the cognitive processes appropriate for instruction and assessment. At the most granular level of the hierarchy, subtopics have seven attributes: (i) prerequisites, (ii) mathematical formulations, (iii) common misconceptions, (iv) engineering applications, (v) cross-cutting topics, (vi) analogies, and (vii) alignment with India’s National Council of Educational Research and Training (NCERT) curriculum (presented as linked references to pertinent sections in NCERT textbooks for grades 8-12).
[…]
@@
The content was reviewed by physics subject matter experts and iteratively refined to ensure accuracy and pedagogical appropriateness. In future iterations, we plan to align with the Next Generation Science Standards (NGSS) [36]. This comprehensive hierarchical organization of concept maps provides the foundation for our automated MCQ generation system to create contextually and pedagogically aligned MCQs. Beyond question generation, this concept map represents a valuable one-time investment with multiple educational applications. The structured knowledge enables intelligent tutoring systems by mapping learning pathways based on prerequisites, supports targeted remediation using documented misconceptions, informs curriculum design, and enhances educational analytics. Structured knowledge also facilitates cross-disciplinary connections by explicitly linking physics concepts to other STEM subjects.
@@ 2.3.2 Concept Map-Based MCQ Pipeline
Figure 1(2) illustrates our end-to-end pipeline for concept map-based MCQ generation, consisting of three main components: (a) prompting with structured content, (b) MCQ generation, and (c) validation.
@@
To evaluate our framework and create a benchmark dataset, we generated questions across 50 topics from the OpenStax textbook using all three approaches. For each topic, we generated five questions that target five levels of cognitive ability, resulting in a dataset of 750 MCQs. This dataset, which we call the ‘OpenStax High School Physics MCQ’ dataset, follows our specified JSON format and will be made publicly available to support future research. The prompts, code, and data are available in the technical appendix.
@@ 3.1 Method
Our rubric intentionally addressed specific quality dimensions. The ‘Relevance’ criterion ensured questions targeted the specified physics topic. ‘Correctness’ verified the scientific accuracy of both the question formulation and the designated correct answer. ‘GradeLevel’ assessment examined whether vocabulary, sentence structure, and concept complexity matched the educational stage of the target audience. ‘Similarity’ checks prevented redundancy, ensuring that questions on the same topic approached the concept from different angles. The ‘BloomsLevel’ rating identified which cognitive skill was being tested, allowing us to evaluate if the questions adequately addressed their intended taxonomic level.
@@ 3.2 Analysis
In distractor-focused criteria, concept map-based generation consistently demonstrated superior performance across all metrics: achieving 80% in ‘Plausibility’, nearly 83% in addressing common ‘Misconceptions’, and approximately 84% in maintaining ‘Independence’ between options. In contrast, the baseline RAG and LLM approaches performed notably lower, with the RAG-based generation scoring slightly better (66-68% in all metrics) than the base LLM approach (61-68% range). Specifically, the base LLM approach showed its weakest performance in ‘Plausibility’ at around 61%, while performing marginally better in ‘Independence’ at 68%. RAG approach maintained a more consistent performance across these criteria, with scores hovering around 66-68%, but still significantly behind the concept of map-based generation. These results highlight the particular strength of our concept map-based method in generating pedagogically sound distractors that are not only plausible enough to challenge students but also effectively address common misconceptions while maintaining clear distinctions between options.
@@ 4.1 Method
Guessing Success $\text{Guessing Success Rate}=\frac{\sum_{i=1}^{N}(G_{i}\times C_{i})}{\sum_{i=1}^{N}G_{i}}\times 100$ (2)
Difficulty Weighted $\text{Difficulty Weighted Accuracy}=\frac{\sum_{i=1}^{N}(C_{i}\times D_{i})}{\sum_{i=1}^{N}D_{i}}\times 100,$ (3)
@@ 4.2 Analysis
This interpretation is strongly supported by the analysis of Guessing Success Rate, where the concept map-based generation showed the lowest rate (28.27%), compared to RAG (33.09%) and the base LLM approach (37.10%). This significantly lower Guessing Success Rate indicates that the questions generated using the concept map-based approach were more resistant to successful guessing strategies. When students were uncertain and resorted to guessing, they were less likely to choose the correct answer for concept map-generated questions than for questions created by the other two methods.
@@ 5 Discussion
Our evaluation revealed that structured domain knowledge significantly improves MCQ quality, particularly in terms of pedagogical and distractor effectiveness. Our method significantly outperformed the baseline methods, with three out of four questions meeting all quality criteria, whereas the baseline methods achieved just over one third. The concept map-based generation excelled in expert evaluation metrics compared to LLM and RAG approaches. Statistically significant differences were found in learner-centric evaluations for guessing success rates. Our concept map-based MCQs had lower guessing success compared to the base LLM, suggesting that they were more challenging, less prone to random guessing, and better at assessing deep conceptual understanding rather than surface knowledge.

### v2

@@ 1 Introduction
Multiple choice questions (MCQs), a widely used objective assessment format, require learners to identify the correct answer from options that include plausible but incorrect alternatives, known as distractors [8]. MCQ-based testing serves two functions: assessment and learning enhancement. Studies have shown that the act of retrieving information while answering MCQs strengthens memory retention and deepens understanding in learners [25, 9]. Evidence from empirical studies demonstrates that MCQs accelerate learning outcomes. Their objective nature allows straightforward evaluation and rapid feedback, essential for continuous assessments in a lesson plan [39, 6]. When distractors are thoughtfully designed to mirror common cognitive errors, MCQs can uncover learning deficiencies and reduce random guessing [18]. This allows educators to modify teaching methods and offer targeted student support [11]. However, creating quality MCQs with effective distractors that challenge learners at different cognitive levels is resource intensive and requires significant domain expertise [1]. To bridge this gap, we propose a novel framework that utilizes concept maps and instruction-tuned LLMs for automated MCQ generation and assessment. The GitHub repository for the code, concept map, and data is available as a supplementary resource for researchers interested in replicating or building upon our work. [^3] GitHub Repository.
@@
Traditional MCQ generation tackled question stem generation, key answer identification, and distractor generation separately. The generation of question stems used rule-based transformations and templates, the identification of key answers relied on domain-specific keywords and phrase matching [12], and corpus-based methods focusing on linguistic features or knowledge-based methods for distractors [23]. These methods suffered from rigid question templates, a lack of contextual understanding in key answer identification, and difficulties in generating plausible semantically relevant distractors, creating major bottlenecks in the creation of high-quality MCQs at scale [1]. Deep learning-based approaches, particularly sequence-to-sequence models [37] with attention mechanisms [4], offered improvements, particularly the hierarchical recurrent encoder-decoder (HRED) architecture [17] with dynamic and static attention [22]. Subsequent studies explored different attention mechanisms to enhance distractor generation, such as co-attention further [33] for passage-question interaction [47], topic-aware attention using Latent Dirichlet Allocation [34], and SoftSel to mitigate answer-revealing distractors [41].
@@
Large-scale datasets played a crucial role in improving MCQ and distractor generation. While datasets like RACE [44] and SWAG [44] focused on assessing reading comprehension and reasoning abilities, domain-specific datasets such as SciQ [43] and EduQG [19] improved structured question generation in different subjects. Distractor selection based on semantic similarity has been done employing cosine similarity of word vectors using these datasets, pre-trained language models (PLMs) such as word2vec [28], GloVe [29], and fastText [7]. Transformer-based models, such as GPT [30], BERT [13], T5 [31], and BART [21], further enhanced distractor generation through fine-tuning and retrieval-augmented pre-training approaches. Recent studies leverage in-context learning [27] and single-stage and multistage prompting techniques [14, 24] to improve distractor quality. These investigations have identified significant room for improvement in the generation of domain-specific and cognitively diverse questions with effective distractors. Furthermore, while promising, LLMs introduce new challenges such as bias [16] and hallucination [45], which could lead to seemingly accurate, yet factually inaccurate key answers and distractors. Although knowledge integration shows promise in grounding model output, current methods still struggle to fully address these challenges, necessitating expert validation, which creates significant resource bottlenecks.
@@ 2.1 Content and Concept Map Design
The concept map is the central piece of our automated MCQ generation framework. To systematically organize the content, we created a hierarchical concept map encompassing 19 units of classical physics using the OpenStax High School Physics textbook [40]. These units encompassed five major topics in physics (Figure 1(1a)): Mechanics, Electricity and Magnetism, Waves and Acoustics, Thermodynamics, and Optics. This hierarchical framework represents a structured knowledge repository to accommodate future expansion with additional subjects and grade levels.
@@
Within our concept map (Figure 1(1b)), each ‘unit’ is aligned with a unit from the OpenStax textbook, encapsulating ‘topics’ that are further divided into ‘subtopics’. ‘Topics’ align with sections in the OpenStax textbook, and for each topic, we mapped learning objectives to the cognitive and knowledge dimensions of the revised Bloom’s taxonomy [3]. This mapping ensures that the knowledge structure not only captures content but also the cognitive processes appropriate for instruction and assessment. We consider question difficulty through cognitive levels of Bloom’s taxonomy, where higher-order thinking skills are considered more difficult than lower-order skills. At the most granular level of the hierarchy, subtopics have seven attributes: (i) prerequisites, (ii) mathematical formulations, (iii) common misconceptions, (iv) engineering applications, (v) cross-cutting topics, (vi) analogies, and (vii) alignment with India’s National Council of Educational Research and Training (NCERT) curriculum (presented as linked references to pertinent sections in NCERT textbooks for grades 8-12).
The prerequisites are divided into physics-specific concepts (e.g., understanding velocity before acceleration) and external knowledge from other subjects (such as mathematical skills), allowing the system to account for knowledge dependencies. The subtopics contain detailed mathematical formulations corresponding to the concept. For example, an equation such as the one for kinetic energy is represented as a JSON object that includes the equation itself formatted using LaTeX syntax (e.g., \$KE = $\frac${1}{2} m vˆ2\$), which allows for accurate mathematical rendering upon retrieval. This data object also contains definitions for each variable and the SI units involved in the equation. Common misconceptions identified through literature and teaching experience inform the creation of targeted assessment questions. Cross-cutting topics highlight connections between physics concepts and other disciplines, enabling interdisciplinary assessment. Engineering applications provide a real-world context that demonstrates the practical relevance of theoretical concepts. Carefully selected analogies help bridge abstract physics concepts with familiar experiences.
@@
The content was reviewed by physics subject matter experts and iteratively refined to ensure accuracy and pedagogical appropriateness. Building the physics concept map required approximately 80 expert-hours (with 4 subject-matter experts each spending 20 hours), including creation and validation of the concept map. We experimented with creating the concept map in a semi-automated way using larger and expensive LLM APIs with in-context examples and found that this could reduce expert effort, though expert validation remains essential. While our current work focuses on physics, the hierarchical framework with grades, subjects, units, topics, and subtopics with detailed information is domain-agnostic. In future iterations, we plan to align with the Next Generation Science Standards (NGSS) [36]. This comprehensive hierarchical organization of concept maps provides the foundation for our automated MCQ generation system to create contextually and pedagogically aligned MCQs. Beyond question generation, this concept map represents a valuable one-time investment with multiple educational applications. The structured knowledge enables intelligent tutoring systems by mapping learning pathways based on prerequisites, supports targeted remediation using documented misconceptions, informs curriculum design, and enhances educational analytics. Structured knowledge also facilitates cross-disciplinary connections by explicitly linking physics concepts to other STEM subjects.
@@ 2.3 MCQ Generation Methods
Algorithm 1 Concept Map-Based MCQ Generation Pipeline
1:User input $u$, Grade level $g$, Cognitive skills $\mathcal{S}=\{\text{Remember, Understand, Apply, Analyze, Evaluate}\}$
2:Assessment with validated MCQs $\mathcal{Q}=\{q_{1},q_{2},...,q_{|\mathcal{S}|}\}$ in JSON format
[…]
9:for each $s\in\mathcal{S}$ do $\triangleright$ Generate questions for each cognitive skill level
10:  $q\leftarrow\textsc{GenerateValidQuestion}(s,t_{user},context,\mathcal{Q})$ $\triangleright$ Generate valid question for skill level
11:  if $q\neq\text{NULL}$ then
12:   $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{q\}$ $\triangleright$ Add valid question to assessment
13:  end if
14:end for
15:return $\mathcal{Q}$
16:function GenerateValidQuestion($s,t,context,\mathcal{H}$)
17:  $max\_attempts\leftarrow 2$ $\triangleright$ Prevent infinite loops
18:  for $attempt=1$ to max_attemptsmax\_attempts do
19:   $question\_history\leftarrow\{q.\text{question}:q\in\mathcal{H}\}$ $\triangleright$ Extract previous questions
20:   $prompt\leftarrow\textsc{BuildPrompt}(context,s,question\_history)$ $\triangleright$ Construct generation prompt
21:   $q\leftarrow\textsc{LLMGenerate}(prompt,\text{temperature}=0.75,\text{CoT}=\textbf{true})$ $\triangleright$ Generate MCQ with Chain-of-Thought
22:   if ValidateJSONFormat(qqthen $\triangleright$ Check JSON format compliance
23:     $eval\leftarrow\textsc{EvaluateQuestion}(q,question\_history)$ $\triangleright$ Detailed evaluation
24:     if $eval.\text{valid}$ then
25:      return $q$ $\triangleright$ Question passes all checks
26:     else
27:      $q_{fixed}\leftarrow\textsc{FixQuestion}(q,eval,question\_history)$ $\triangleright$ Attempt to fix issues
28:      $eval_{fixed}\leftarrow\textsc{EvaluateQuestion}(q_{fixed},question\_history)$
29:      if $eval_{fixed}.\text{valid}$ then
30:        return $q_{fixed}$
31:      end if
32:     end if
33:   end if
34:  end for
35:  return NULL $\triangleright$ Failed to generate valid question
36:end function
37:function EvaluateQuestion($q,\mathcal{H}$)
38:  Evaluate uniqueness: $unique\leftarrow\textsc{CheckUniqueness}(q,\mathcal{H})$
39:  Evaluate correctness: $correct\leftarrow\textsc{CheckAnswerCorrectness}(q)$
40:  $valid\leftarrow unique\land correct$
41:  return $\{valid,unique,correct,issues\}$
42:end function
43:function FixQuestion($q,eval,\mathcal{H}$)
44:  if $\neg eval.unique$ then
45:   return GenerateNewQuestion($q.skill$, $\mathcal{H}$) $\triangleright$ Generate completely new question
46:  else if $\neg eval.correct$ then
47:   return FixAnswerOnly($q$) $\triangleright$ Fix only answer and explanation
48:  end if
49:  return $q$
50:end function
@@ 2.3.2 Concept Map-Based MCQ Pipeline
Figure 1(2) illustrates our end-to-end pipeline for concept map-based MCQ generation, consisting of three main components: (a) prompting with structured content, (b) MCQ generation, and (c) validation. The pipeline is also detailed as an algorithm in Algorithm 1.
@@
To evaluate our framework and create a benchmark dataset, we generated questions across 50 topics from the OpenStax textbook using all three approaches. For each topic, we generated five questions that target five levels of cognitive ability, resulting in a dataset of 750 MCQs. This dataset, which we call the OpenStax PhyQ dataset, follows our specified JSON format and is publicly available to support future research in the GitHub Repository [^6] GitHub Repository.
@@ 3.1 Method
Agreement 1.00 1.00 0.99 0.96 0.90 0.96 0.98 0.99
Kappa 1.00 1.00 0.83 0.86 0.86 0.62 0.56 0.58
Our rubric intentionally addressed specific quality dimensions. The ‘Relevance criterion ensured that the questions targeted the specified physics topic. ‘Correctness’ verified the scientific accuracy of both the question formulation and the designated correct answer. ‘GradeLevel’ assessment examined whether vocabulary, sentence structure, and concept complexity matched the educational stage of the target audience. ‘Similarity’ checks prevented redundancy, ensuring that questions on the same topic approached the concept from different angles. The ‘BloomsLevel’ rating identified which cognitive skill was being tested, allowing us to evaluate if the questions adequately addressed their intended taxonomic level.
@@ 3.2 Analysis
In distractor-focused criteria, concept map-based generation consistently demonstrated superior performance across all metrics: achieving 80% in ‘Plausibility’, nearly 83% in addressing common ‘Misconceptions’, and approximately 84% in maintaining ‘Independence’ between options. In contrast, the baseline RAG and LLM approaches performed notably lower, with the RAG-based generation scoring slightly better (66-68% in all metrics) than the base LLM approach (61-68% range). Specifically, the base LLM approach showed its weakest performance in ‘Plausibility’ at around 61%, while performing marginally better in ‘Independence’ at 68%. RAG approach maintained a more consistent performance across these criteria, with scores hovering around 66-68%, but still significantly behind the concept map-based generation. These results highlight the particular strength of our concept map-based method in generating pedagogically sound distractors that are not only plausible enough to challenge students but also effectively address common misconceptions while maintaining clear distinctions between options.
@@ 4.1 Method
$\text{Guessing Success Rate}=\frac{\sum_{i=1}^{N}(G_{i}\times C_{i})}{\sum_{i=1}^{N}G_{i}}\times 100$ (2)
$\text{Difficulty Weighted Accuracy}=\frac{\sum_{i=1}^{N}(C_{i}\times D_{i})}{\sum_{i=1}^{N}D_{i}}\times 100,$ (3)
@@ 4.2 Analysis
This interpretation is strongly supported by the analysis of Guessing Success Rate, where the concept map-based generation showed the lowest rate (28.05%), compared to RAG (33.16%) and the base LLM approach (37.10%). This significantly lower Guessing Success Rate indicates that the questions generated using the concept map-based approach were more resistant to successful guessing strategies. When students were uncertain and resorted to guessing, they were less likely to choose the correct answer for concept map-generated questions than for questions created by the other two methods.
@@ 5 Discussion
Our evaluation revealed that structured domain knowledge significantly improves MCQ quality, particularly in terms of pedagogical and distractor effectiveness. Our method significantly outperformed the baseline methods, with three out of four questions meeting all quality criteria, whereas the baseline methods achieved just over one third. The concept map-based generation excelled in expert evaluation metrics compared to LLM and RAG approaches. Statistically significant differences were found in learner-centric evaluations for guessing success rates. Our concept map-based MCQs had lower guessing success compared to the base LLM, suggesting that they were more challenging, less prone to random guessing, and better at assessing deep conceptual understanding rather than surface knowledge. These results underscore the fundamental advantages of structured knowledge representation in educational content generation. While stronger LLMs might improve baselines, structured grounding through concept maps provides distinct advantages that extend beyond model capabilities. Concept map creation represents a one-time investment that enables consistent, pedagogically sound generation—a process that could be semi-automated with a capable LLM and some in-context examples. Without such a structured foundation, even expensive models require extensive prompting and often produce distractors that could increase guessing success rates, undermining the assessment’s effectiveness.
@@ 6 Conclusion
{ack}
We are grateful to the administration, faculty, and staff of Kendriya Vidyalaya, Indian Institute of Science, Bengaluru, for their institutional support and cooperation in facilitating this research. We sincerely thank all students, whose participation was essential for the successful completion of this study.